OpenStudy (anonymous):

Factor completely: x2 − 3x + 5

OpenStudy (anonymous):

(x+a)(x+b) = x^2 + (a+b)x + ab thus... a+b = -3 ab = 5 b = 5/a a + (5/a) = -3 plug it in your graphing calculator, should give you a, then plug that value into ab = 5 to get b.